OpenStudy (anonymous):

Convert the following decimal negative numbers to their 8-bit, 2’s complement formulation: -1, -99, -123

OpenStudy (anonymous):

do you know how to find the 1's complement for the numbers? the 2's complement = 1's complement + 1 or you can use the general formula, the n-bit 2's complement of a positive integer X, \[X'=2^{n}-X\]

OpenStudy (anonymous):

i'll do 1 example. +1 in 8-bit representation = 0000 0001 -1 in 1's complement = 1111 1110 -1 in 2's complement = 1111 1111 using the formula, \[X'=2^8-1\]\[X'=256-1=255_{10}=11111111_{2}\]